How many arrangement are there of the letters DISAPPOINT?


1
Expert's answer
2021-03-24T14:06:03-0400

10 is the total number of letters where 2 times P, 2 times I occurs in DISAPPOINT.

"\frac{10!}{2!2!} = \frac{2 \times 3 \times 4 \times 5 \times 6 \times 7 \times 8 \times 9 \times 10}{2 \times 2} = 907200"